2. The Blue Company makes boats. Job 589 is to build 10 boats. The company has 2 departments – assembly and fabrication. The assembly department allocates manufacturing overhead to jobs based on direct labor hours. The fabrication department allocates manufacturing overhead to jobs based on machine hours. Please use the following information about the Blue Company; Assembly Department Fabrication Department Total fixed manufacturing overhead cost estimated for all jobs during the whole year $480,000 $560,000 Estimated direct labor hours for all jobs during the whole year 80,000 60,000 Estimated machine hours for all jobs during the whole year 20,000 70,000 Estimated variable manufacturing overhead cost per direct labor hour $9 ------ Estimated variable manufacturing overhead cost per machine hour ----- $6 The following data is for Job 589: Assembly Department Fabrication Department Direct materials cost $400 $600 Direct labor cost $900 $300 Direct labor hours used 70 20 Machine hours used 30 50 Please calculate for Job 589: a. the total direct materials cost b. the total direct labor cost c. The predetermined overhead rate for the Assembly department d. The predetermined overhead rate for the Fabrication department e. The manufacturing overhead cost applied to job 589 by the Assembly department f. The manufacturing overhead cost applied to job 589 by the Fabrication department g. The total manufacturing overhead cost applied to job 589 h. The total cost of job 589 i. The cost per unit (per boat)
Answer #1

Solution

Blue Company

  1. Calculation of total material cost for job 589:

Total material cost = $400 + $600 = $1,000

  1. Total direct labor cost = $900 + $300 = $1,200
  2. Predetermined overhead rate for the assembly department –

Predetermined overhead rate = total fixed manufacturing overhead/direct labor hours

Fixed manufacturing overhead for Assembly department = $480,000

Direct labor hours = 80,000

Predetermined overhead rate = 480,000/80,000 = $6 per direct labor hour

  1. The predetermined overhead rate for the fabrication department:

Predetermined overhead rate = total fixed manufacturing overhead/machine hours

Fixed manufacturing overhead for Fabrication department = $560,000

Machine hours = 70,000

Predetermined overhead rate = 560,000/70,000 = $8 per machine hour

  1. Manufacturing overhead cost applied to job 589 by the assembly department:

Manufacturing overhead rate applied to job 589 = direct labor hours x predetermined overhead rate at assembly department

Direct labor hours = 70

Predetermined overhead rate = $6

Manufacturing overhead applied = $6 x 70 = $420

  1. Manufacturing overhead cost applied to job 589 by the fabrication department:

Manufacturing overhead rate applied to job 589 = machine hours x predetermined overhead rate at fabrication department

machine hours = 50

Predetermined overhead rate = $8

Manufacturing overhead applied = $8 x 50 = $400

  1. Total manufacturing overhead applied to job 589 –

Total manufacturing overhead applied to Job 589 = Manufacturing overhead cost applied to job 589 by the assembly department + Manufacturing overhead cost applied to job 589 by the fabrication department

= $420 + $400 = $820

  1. Total cost of job 589 –

Direct material cost = $1,000

Direct labor cost s= $1,200

Variable MOH cost = $1,290

Fixed Manufacturing overhead = $820

Total cost = $4,310

Note: computation of variable manufacturing overhead –

Variable manufacturing overhead = total direct labor hours x variable overhead rate per DLH

Direct labor hours = 70 + 20 = 90

Variable MOH per DLH = $9

Variable MOH = $9 x 90 = $810

Variable manufacutirng overhead cost for machine hours –

Machine hours used = 30 + 50 = 80

Variable MOH per machine hour = $6

Variable MOH = $6 x 80 = $480

Total variable MOH for Job589 = $810 + $480 = $1,290

  1. The cost per unit (per boat):

Cost per boat = total cost/number of boats in Job 589

Total cost = $4,310

Number of boats = 10

Cost per boat = $4,310/10 = $431

Hence, cost per boat = $431